# Topological potentials guiding protein self-assembly

### v2

@@ 2.3.4 Analyzing the impact of the topological potential in different simulation phases
$F_{sol}^{*}$ finish, $T=3.0K$ $1\%$ $6\%$
$E_{comb}$ finish, $T=3.0K$ $25\%$ $63\%$
$E_{comb}$ finish, $T=5.0K$ $36\%$ $54\%$
@@
We simulate $E_{comb}$ for both $T=3.0K$ and $T=5.0K$ and see that in both cases simulations are more likely to succeed on the set of states that were aligned by $E_{comb}$. Interestingly, the lower temperature performs better in the $E_{comb}$ aligned states. We assume that this is caused by the selected states being comparatively good and that a higher temperature yields a higher chance to move away from them than a lower temperature. What is a drawback for well-aligned states is a feature for the low-quality alignments that appear during pure $F_{sol}^{*}$ simulations. Here, the higher temperature performs better, because moving away from the initialized alignment and realignment driven by $E_{comb}$ is favorable.
@@ 2.3.5 Improvements for other systems of protein assembly
To show that this approach can be generalized, we apply it to simulate dimerization of the human hepatitis B virus core protein Alexander et al. (2013), dimerization of an accessory protein of the SARS-CoV-2 virus that hinders the immune response Weeks et al. (2020), and trimerization of an extracellular fragment of the human CD40 ligand, relevant to T cell function Karpusas et al. (1995a). All of these structures are assemblies of homodimers and homotrimers, respectively, which means that they are composed of two or three identical protein chains. Figure 6 shows the renderings of the final assembled states found via simulations driven by
@@
We ran simulations for the pure morphometric approach and for combinations of $F_{sol}^{*}$ with $\mathcal{T}_{(1.0,-0.4,-0.9)}$ at $\mu=0.3$ and $\mathcal{T}_{(1.0,-1.0,1.0)}$ at $\mu=0.8$, corresponding to the parameters with the best performance, for the respective $\lambda_{i}$, observed in Figure 4. We ran $40$ simulations each at simulation temperatures $T\in\{2.0K,3.0K,4.0K,5.0K\}$ for two subunits and at $T=5.0K$ for three subunits.
@@ Acknowledgments
The geometric and topological measures are derived from the weighted Alpha complex of the atomic centers. The Alpha complex is a subcomplex of the weighted Delaunay triangulation that provides a mathematically precise and computationally efficient representation of the union-of-balls model of a molecule Edelsbrunner and Mücke [1994], Edelsbrunner [1992]. For a set of atomic centers $P=\{p_{1},\dots,p_{n}\}\subset\mathbb{R}^{d}$ and a corresponding set of weights $W=\{w_{1},\dots,w_{n}\}\subset\mathbb{R}$, the construction begins with the weighted Voronoi diagram. The $i$-th weighted Voronoi cell is defined as
@@
i.e. the dual of the weighted Voronoi cells intersected with the balls $B(p_{i},\sqrt{w_{i}})$. Inclusion-Exclusion formulas to compute the geometric measures $V$, AA, $C$ and $X$, are based on this incidence information Edelsbrunner [1993], Edelsbrunner and Koehl [2003], Bryant et al. [2004], Akopyan and Edelsbrunner [2020a, b]. These formulas are implemented in the AlphaMol program Koehl [2021]. We also use the weighted Alpha complex to compute the overlap penalty $L$, by iterating over the edges whose vertices correspond to different proteins and summing up the overlap of the corresponding balls.
@@
are the $p$-dimensional persistent Betti numbers. They count the $p$-dimensional cycles that are not boundaries. Given a class $\gamma\in H_{p}(\mathcal{A}_{i})$ we say $\gamma$ is born at $\mathcal{A}_{i}$ if $\gamma\notin H_{p}^{i-1,i}$. We say it dies at $\mathcal{A}_{j}$ if $f_{p}^{i,j-1}(\gamma)\notin H_{p}^{i-1,j-1}$ and $f_{p}^{i,j}(\gamma)\in H_{p}^{i-1,j}$. We call the interval $(\alpha_{i},\alpha_{j})$ the birth-death interval of $\gamma$ and $\alpha_{j}-\alpha_{i}$ its persistence. We refer to the collection of all $p$-dimensional birth-death intervals as the $p$-dimensional persistence diagram and denote it by $Dgm_{p}$ Edelsbrunner and Harer [2010].
We use the Oineus software package Nigmetov et al. [2025] to calculate persistent homology.
@@
We simulate configurations of protein subunits using variants of a random walk Metropolis Hastings [1970]. Each configuration is represented by a state $x\in(\operatorname{SO}_{3}\times\mathbb{R}^{3})^{m}$, where $m$ is the number of simulated molecules. The geometries of the systems we study are taken from the protein data base entries with identifiers 6R7M Schmidli et al. [2019], 4BMG Alexander et al. [2013], 6Z4U Weeks et al. [2020] and 1ALY Karpusas et al. [1995b], for TMV, hepatitis B, SARS-CoV2 and the CD40 ligand respectively. For atomic radii we use the ProtOr radii Tsai et al. [1999], which account for hydrogen atoms by giving radii for atomic groups instead of individual atoms. Each simulation is run in a bounding box of size $120\text{\AA }$. These bounds are not checked for each individual atom but for each center of mass of the simulated molecules.
@@
Assuming the current state of the simulation is state $x^{i}$, we pick one of the simulated molecules at random and perturb it to generate a candidate state $x^{i+1}$. This is also known as component-wise or block-wise random walk Metropolis Johnson et al. [2009].
The probability of accepting $x^{i+1}$ as the new state of the simulation is determined as:
@@
where $\Delta E=E(x^{i+1})-E(x^{i})$ is the difference in potential energy of the states. $T$ is a simulation temperature, which regulates the probability of accepting states with higher potential energy. The potential energy used here is $E_{comb}$ as in Equation 3 for different parameters $\mu$ and $\lambda_{i}$. The prefactors of $F_{sol}^{*}$ are given by the White Bear prefactors Hansen-Goos and Roth [2006] computed for packing fraction $\eta=0.3665$ and solvent radius $r_{s}=1.4\text{\AA }$, mimicking the properties of water. For the simulated annealing (SA) Kirkpatrick et al. [1983] simulations we change $T$ over the course of a simulation by letting it decay to zero according to a quadratic annealing schedule:
@@
where $I$ is the total number of iterations at which the temperature reaches zero and $k$ is the current step in the decay period. Once $T=0.0K$ is reached, we reheat the system several times to scan for nearby local minima.
@@
The morphometric approach is short ranged and additive. As such it is not necessary to recompute the potential energy of the whole system in every step, but only the connected components of molecules that have changed. We determine the connected components in each step by constructing a graph that has each molecule as a node and an edge whenever the bounding spheres of two molecules overlap. Let idxidx be the index of the molecule we moved for the proposal of simulation step $i+1$. Let $\mathcal{C}_{i}$ be the set of connected components from step $i$ and let $\mathcal{C}_{i+1}$ be the set of connected components from step $i+1$. For each component $c\in\mathcal{C}_{i+1}$ there are three cases to consider. Case 1: The connected component $c$ is an element of $\mathcal{C}_{i}$ and $idx\notin c$. Then the potential energy of $c$ remains unchanged and we can just copy the energy value from the last iteration. Case 2: The connected component $c$ is an element of $\mathcal{C}_{i}$ and $idx\in c$. Then the potential energy of $c$ changes and we have to recompute it. Case 3: The connected component $c$ is not an element of $\mathcal{C}_{i}$. Then the potential energy of $c$ is changed and we need to recompute. By passing the potential energy value of a single molecule to the algorithm we can further optimize this procedure and not recalculate the potential energy of components that consists of only one molecule.
To ensure a fair comparison between different topological potential setups, we defined a relative contribution measurement, $c_{\mathcal{T}}$. This measurement quantifies the energetic impact of the topological term relative to the geometric term during the last stage of the assembly process. We take the set of $800$ simulations, the data shown in Figure 4 A are based on. For every simulation $i$ we take two states $S^{i}_{min}$, the minimal potential energy state, and $S^{i}_{disp}$, the state immediately before the solvent-accessible surface areas of the protein subunits start to interact, before $S^{i}_{min}$ is reached. We then calculate an average of how much of the change in potential energy in the final phase in these simulations is contributed by $\mathcal{T}$ in relation to $F_{sol}^{*}$ as:
@@
for the solvent radius $r_{s}$ are the weights of the weighted Alpha complex used to compute the morphometric measures. Now let $\mathcal{B}(\alpha)=\bigcup\limits_{j=1}^{m}B_{j}(\alpha)$ where $B_{j}(\alpha)$ is the ball centered at $p_{j}$ with radius $\sqrt{w_{j}+\alpha}$. If $w_{j}+\alpha<0$, we define $B_{j}(\alpha)=\emptyset$. Because the Alpha complex $\mathcal{A}_{\alpha}$ is homotopy equivalent to the union of balls $\mathcal{B}(\alpha)$, their Euler characteristics are equal Edelsbrunner [1993]. The sequence of values, $\alpha_{0}<\dots<\alpha_{k}$ for which topological changes occur, when moving from $\mathcal{A}(\alpha_{i})$ to $\mathcal{A}(\alpha_{i+1})$, are the filtration values of the Alpha filtration on which we calculate persistent homology. Here
@@
Then by the Euler-Poincaré formula Edelsbrunner and Harer [2010] we get:
@@
This problem can be solved efficiently using the Kabsch algorithm Kabsch [1978].
@@
The code we developed to generate the data presented in this study is available on github Spirandelli [2025].

### v3

@@ 2.3.4 Analyzing the impact of the topological potential in different simulation phases
$F_{sol}^{*}$ finish, $T=3.0$ $1\%$ $6\%$
$E_{comb}$ finish, $T=3.0$ $25\%$ $63\%$
$E_{comb}$ finish, $T=5.0$ $36\%$ $54\%$
@@
We simulate $E_{comb}$ for both $T=3.0$ and $T=5.0$ and see that in both cases simulations are more likely to succeed on the set of states that were aligned by $E_{comb}$. Interestingly, the lower temperature performs better in the $E_{comb}$ aligned states. We assume that this is caused by the selected states being comparatively good and that a higher temperature yields a higher chance to move away from them than a lower temperature. What is a drawback for well-aligned states is a feature for the low-quality alignments that appear during pure $F_{sol}^{*}$ simulations. Here, the higher temperature performs better, because moving away from the initialized alignment and realignment driven by $E_{comb}$ is favorable.
@@ 2.3.5 Improvements for other systems of protein assembly
To show that this approach can be generalized, we apply it to simulate dimerization of the human hepatitis B virus core protein [35], dimerization of an accessory protein of the SARS-CoV-2 virus that hinders the immune response [36], and trimerization of an extracellular fragment of the human CD40 ligand, relevant to T cell function [37]. All of these structures are assemblies of homodimers and homotrimers, respectively, which means that they are composed of two or three identical protein chains. Figure 6 shows the renderings of the final assembled states found via simulations driven by
@@
We ran simulations for the pure morphometric approach and for combinations of $F_{sol}^{*}$ with $\mathcal{T}_{(1.0,-0.4,-0.9)}$ at $\mu=0.3$ and $\mathcal{T}_{(1.0,-1.0,1.0)}$ at $\mu=0.8$, corresponding to the parameters with the best performance, for the respective $\lambda_{i}$, observed in Figure 4. We ran $40$ simulations each at simulation temperatures $T\in\{2.0,3.0,4.0,5.0\}$ for two subunits and at $T=5.0$ for three subunits.
@@ Acknowledgments
The geometric and topological measures are derived from the weighted Alpha complex of the atomic centers. The Alpha complex is a subcomplex of the weighted Delaunay triangulation that provides a mathematically precise and computationally efficient representation of the union-of-balls model of a molecule [28, 29]. For a set of atomic centers $P=\{p_{1},\dots,p_{n}\}\subset\mathbb{R}^{d}$ and a corresponding set of weights $W=\{w_{1},\dots,w_{n}\}\subset\mathbb{R}$, the construction begins with the weighted Voronoi diagram. The $i$-th weighted Voronoi cell is defined as
@@
i.e. the dual of the weighted Voronoi cells intersected with the balls $B(p_{i},\sqrt{w_{i}})$. Inclusion-Exclusion formulas to compute the geometric measures $V$, AA, $C$ and $X$, are based on this incidence information [38, 39, 40, 41, 42]. These formulas are implemented in the AlphaMol program [31]. We also use the weighted Alpha complex to compute the overlap penalty $L$, by iterating over the edges whose vertices correspond to different proteins and summing up the overlap of the corresponding balls.
@@
are the $p$-dimensional persistent Betti numbers. They count the $p$-dimensional cycles that are not boundaries. Given a class $\gamma\in H_{p}(\mathcal{A}_{i})$ we say $\gamma$ is born at $\mathcal{A}_{i}$ if $\gamma\notin H_{p}^{i-1,i}$. We say it dies at $\mathcal{A}_{j}$ if $f_{p}^{i,j-1}(\gamma)\notin H_{p}^{i-1,j-1}$ and $f_{p}^{i,j}(\gamma)\in H_{p}^{i-1,j}$. We call the interval $(\alpha_{i},\alpha_{j})$ the birth-death interval of $\gamma$ and $\alpha_{j}-\alpha_{i}$ its persistence. We refer to the collection of all $p$-dimensional birth-death intervals as the $p$-dimensional persistence diagram and denote it by $Dgm_{p}$ [43].
We use the Oineus software package [44] to calculate persistent homology.
@@
We simulate configurations of protein subunits using variants of a Random Walk Metropolis [45]. Each configuration is represented by a state $x\in(\operatorname{SO}_{3}\times\mathbb{R}^{3})^{m}$, where $m$ is the number of simulated molecules. The geometries of the systems we study are taken from the protein data base entries with identifiers 6R7M [32], 4BMG [35], 6Z4U [36] and 1ALY [37], for TMV, hepatitis B, SARS-CoV2 and the CD40 ligand respectively. For atomic radii we use the ProtOr radii [33], which account for hydrogen atoms by giving radii for atomic groups instead of individual atoms. Each simulation is run in a bounding box of size $120\text{\AA }$. These bounds are not checked for each individual atom but for each center of mass of the simulated molecules.
@@
Assuming the current state of the simulation is state $x^{i}$, we pick one of the simulated molecules at random and perturb it to generate a candidate state $x^{\prime}$. This is also known as component-wise or block-wise Random Walk Metropolis [46].
The probability of accepting $x^{\prime}$ as the new state of the simulation is determined as:
@@
where $\Delta E=E(x^{\prime})-E(x^{i})$ is the difference in energy of the states. If $x^{\prime}$ is accepted, we set $x^{i+1}=x^{\prime}$. Otherwise we set $x^{i+1}=x^{i}$. $T$ is a simulation temperature, which regulates the probability of accepting states with higher energy. The energy used here is $E_{comb}$ as in Equation 3 for different parameters $\mu$ and $\lambda_{i}$. The prefactors of $F_{sol}^{*}$ are given by the White Bear prefactors [26] computed for packing fraction $\eta=0.3665$ and solvent radius $r_{s}=1.4\text{\AA }$, mimicking the properties of water. For the simulated annealing (SA) [47] simulations we change $T$ over the course of a simulation by letting it decay to zero according to a quadratic annealing schedule:
@@
where $I$ is the total number of iterations at which the temperature reaches zero and $k$ is the current step in the decay period. Once $T=0.0$ is reached, we reheat the system several times to scan for nearby local minima.
@@
The morphometric approach is short ranged and additive. As such it is not necessary to recompute its energetic contribution of the whole system in every step, but only the connected components of molecules that have changed. We determine the connected components in each step by constructing a graph that has each molecule as a node and an edge whenever the bounding spheres of two molecules overlap. Let idxidx be the index of the molecule we moved for the proposal of simulation step $i+1$. Let $\mathcal{C}_{i}$ be the set of connected components from step $i$ and let $\mathcal{C}_{i+1}$ be the set of connected components from step $i+1$. For each component $c\in\mathcal{C}_{i+1}$ there are three cases to consider. Case 1: The connected component $c$ is an element of $\mathcal{C}_{i}$ and $idx\notin c$. Then the energy of $c$ remains unchanged and we can just copy the energy value from the last iteration. Case 2: The connected component $c$ is an element of $\mathcal{C}_{i}$ and $idx\in c$. Then the energy of $c$ changes and we have to recompute it. Case 3: The connected component $c$ is not an element of $\mathcal{C}_{i}$. Then the energy of $c$ is changed and we need to recompute. By passing the energy value of a single molecule to the algorithm we can further optimize this procedure and not recalculate the energetic contributions of components that consists of only one molecule.
To ensure a fair comparison between different topological potential setups, we defined a relative contribution measurement, $c_{\mathcal{T}}$. This measurement quantifies the energetic impact of the topological term relative to the geometric term during the last stage of the assembly process. We take the set of $800$ simulations, the data shown in Figure 4 A are based on. For every simulation $i$ we take two states $S^{i}_{min}$, the minimal energy state, and $S^{i}_{disp}$, the state immediately before the solvent-accessible surface areas of the protein subunits start to interact, before $S^{i}_{min}$ is reached. We then calculate an average of how much of the change in energy in the final phase in these simulations is contributed by $\mathcal{T}$ in relation to $F_{sol}^{*}$ as:
@@
for the solvent radius $r_{s}$ are the weights of the weighted Alpha complex used to compute the morphometric measures. Now let $\mathcal{B}(\alpha)=\bigcup\limits_{j=1}^{m}B_{j}(\alpha)$ where $B_{j}(\alpha)$ is the ball centered at $p_{j}$ with radius $\sqrt{w_{j}+\alpha}$. If $w_{j}+\alpha<0$, we define $B_{j}(\alpha)=\emptyset$. Because the Alpha complex $\mathcal{A}_{\alpha}$ is homotopy equivalent to the union of balls $\mathcal{B}(\alpha)$, their Euler characteristics are equal [38]. The sequence of values, $\alpha_{0}<\dots<\alpha_{k}$ for which topological changes occur, when moving from $\mathcal{A}(\alpha_{i})$ to $\mathcal{A}(\alpha_{i+1})$, are the filtration values of the Alpha filtration on which we calculate persistent homology. Here
@@
Then by the Euler-Poincaré formula [43] we get:
@@
This problem can be solved efficiently using the Kabsch algorithm [48].
@@
The code we developed to generate the data presented in this study is available on github [49].